OpenStudy (anonymous):

H9P1: RESPONSE TO A DELAYED IMPULSE A=2/π=0.64 , t=9.0s. At t=0 the state is: vC(0)=0.0 and iL(0)=1.0. B. At the initial time what is the total energy, in Joules, stored in the circuit? C. At the time just before the impulse happens t=9.0− what is the total energy, in Joules, stored in the circuit? E. At the time just before the impulse happens what is the voltage vC(9.0−), in Volts, across the capacitor?

OpenStudy (anonymous):

B) n C) total energy is ½*L*iL^2 (in Joules) E) vC = -L*2*pi*f*sin(2*pi*f*t)

OpenStudy (anonymous):

Lab 9 Figure 2. Boost converter

OpenStudy (anonymous):

what is the ans. of Lab 9 Figure 2. Boost converter

OpenStudy (anonymous):

double click on square part of d diagram and change the duty cycle % to 58 and change it after that click on TRAN option on top-left of diagram n without closing d dialog box click check :)

OpenStudy (anonymous):

H10P3: An L Network Cmatch=? Lmatch=?

OpenStudy (anonymous):

giv me the value of f for this question pls

OpenStudy (anonymous):

R1=1000.0Ω. R2=50.0Ω f=630.0kHz

OpenStudy (anonymous):

Lmatch - 55.05 Cmatch - 1159.13

OpenStudy (anonymous):

Plzzz Help me. H10P3: An L Network if f=630.0kHz Cmatch=????? Lmatch=?????

OpenStudy (anonymous):

Lmatch - 55.05 is correct bt Cmatch - 1159.13 in incorrect

OpenStudy (anonymous):

?

OpenStudy (anonymous):

w = 2*pi*f Cmatch = R1R2/(w*sqrt(R1R2 – R2^2)) or Cmatch = 1/(w*sqrt(R1R2 – R2^2)) calculate n see wich 1 is correct :)

OpenStudy (anonymous):

pi=?

OpenStudy (anonymous):

3.14

OpenStudy (anonymous):

R1=1000.0Ω. R2=50.0Ω. f=630.0kHz w= 2*pi*f = 3956.4 but according to this formula Cmatch = R1R2/(w*sqrt(R1R2 – R2^2)) = 0.05798598066 Cmatch = 1/(w*sqrt(R1R2 – R2^2)) = 1.159719613*10^-6 both are incorrect

OpenStudy (anonymous):

w=3958406.744 Cmatch = 1/(w*sqrt(R1R2 – R2^2)) = 1159.131685 Cmatch = R1R2/(w*sqrt(R1R2 – R2^2)) =57956584.23 (surely wrong dis one)

OpenStudy (anonymous):

haa yr

OpenStudy (anonymous):

log out from ur edx account and den login again and den paste dis ans in ur Cmatch ans 1159.131685 n den chek

OpenStudy (anonymous):

HOW TO CHECK BOOST CONVERTOR MY ANSWER IS STILL NOT CORRECT

OpenStudy (anonymous):

after converting duty cycle % to 58 do transient analysis of ckt and den directly chek it

OpenStudy (anonymous):

At the time just before the impulse happens what is the voltage vC(9.0−), in Volts, across the capacitor? if L=70H, C=5.79mF, t=9sec , iL(0)=1 plzzz tell me the answer.....

OpenStudy (anonymous):

v= -L*2*pi*f*sin(2*pi*f*t) = -70*2*pi*0.25*sin(2*pi*0.25*9) = -109.95574